OpenStudy (anonymous):

slove this 2x+3y=21 3x-6y=-24 slove this with substution method

OpenStudy (anonymous):

2x + 3y = 21 -3x - 6y = -24 Solve the second equation for x -3x - 6y = -24 -3x = -24 + 6y Divide every term by -3 x = 8 - 2y Substitute (8-2y) for x in the OTHER (first) equation 2x + 3y = 21 2(8-2y) + 3y = 21 16-4y + 3y = 21 16 - y = 21 -y = 5 y = -5 Substitute -5 for y in x = 8-2y x = 8 - 2y x = 8 - 2(-5) x = 8 + 10 x = 18 So the solution is (x, y) = (18, -5)